Consider two events X and Y that have marginal probabilities of 0.68 and 0.57 respectively. Their joint probability is 0.35. Calculate the probability that event X occurs if Y has occurred. (3 decimal places)
Answer #1

Conditional probability is the probability of intersection of X and Y divided by the marginal probability of an event which has already occurred.

Pr(X/Y) = Pr(X,Y)/Pr(Y) =0.35/ 0.57= 0.614